OpenStudy (anonymous):

Simplify. Write in standard form. t2 (t + 1) – t (2t2 – 1) needs the steps and i will give medals

zepdrix (zepdrix):

\(\Large\bf \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf t^2(t+1)-t(2t^2-1)\]

zepdrix (zepdrix):

So you'll start by distributing some stuff.

zepdrix (zepdrix):

\[\Large\bf\sf \color{orangered}{t^2}(t+1)-t(2t^2-1)\quad=\quad t\cdot\color{orangered}{t^2}+1\cdot\color{orangered}{t^2}-t(2t^2-1)\]Which simplifies to,\[\Large\bf\sf t^3+t^2-t(2t^2-1)\]We distributed the t^2 to each term in the first set of brackets. Understand how that works? Think you can do it with the other set of brackets? You have to distribute a -t to each term so it's a little trickier.